OpenStudy (anonymous):

solve: -3/x-2 = 1 - 12/x^2-4

OpenStudy (anonymous):

You want to find common terms here. There's a very common, easy to spot little trick you can use here: x^2 - 4 = (x-2)(x+2) Or in general for any x and any y with an integer root: x^2 - y = (x-y)(x+y) So we can apply that here: -3/(x-2) = (1-12)/(x-2)(x+2) Solve for x now that it's easy.

OpenStudy (anonymous):

how would I go about solving for x? @Rav

OpenStudy (anonymous):

Continuing on: -3/(x-2) = (1-12)/(x-2)(x+2) See on both sides of the equation, you can see (x-2). We want that gone. Multiply both sides by (x-2): -3 = (1-12)/(x+2) Now multiply by (x+2) on both sides: -3(x+2) = 1 - 12 = -11 Negate and simplify: 3x + 6 = 11 Move the 6: 3x = 11 - 6 = 5 Divide by 3: x = 5/3

OpenStudy (anonymous):

im kinda confuse because the only answers it could be is A. -2 B. -5 C. -2,2 or D. 2.-5 @ rav

OpenStudy (anonymous):

@Rav

OpenStudy (anonymous):

Is it 1 - (12)/(x^2 - 4) or is it (1-12)/(x^2-4)?

OpenStudy (anonymous):

its 1- (12(/(x^2-4)

OpenStudy (anonymous):

Fixed: -3/(x-2) = 1 - (12)/(x-2)(x+2) -3 = (x-2) - (12)/(x+2) Multiply by (x+2): -3x - 6 = x^2 - 4 - 12 x^2 - 16 + 3x + 6 = 0 x^2 + 3x - 10 = 0 Now you can use the quadratic formula or you can go further with this method: Now you want (x +/- a)(x +/- b) such that ab = 10 and a + b = 3 Can we have this? Yes: if |a| = 5 and |b| = 3, with one of them negative and one positive. So that's: x^2 + 3x - 10 = (x - 2)(x + 5) = 0 So now you substitute x-values until you find which answer fits.

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

I got the answer B @Rav

OpenStudy (anonymous):

-5 IS a solution to this problem. So is 2. Try out all answers! The correct answer is D.

OpenStudy (anonymous):

ill try it again

OpenStudy (anonymous):

thank you